OpenStudy (anonymous):

Yoki bought x pounds of peanuts for $3.50 per pound and y pounds of cashews for $5.05 per pound at the store for a retirement party. The total bill was $24.10 for 6 pounds of nuts. Part A Write a system of two equations that could be used to find the number of pounds of peanuts and the number of pounds of cashews bought. Part B Solve the system. How many pounds of each nut did Yoki buy?

OpenStudy (radar):

Note that x + y = 6 (total of 6 pound of nuts) Did you catch that. That is one of the equations for this system.

OpenStudy (radar):

That takes care of the pounds. Now follow the money. 3.50x + 5.05y=$24.10 That is your 2nd equation for the system. From 1st express x in terms of y, x=6-y substitute 6-y for x in the 2nd equation and solve.

OpenStudy (radar):

Is this gonna be a monologue?

OpenStudy (anonymous):

Sorry im just trying to figure it out

OpenStudy (radar):

Note the problem assigns what x and y represent and do you see where it tells you that x + y = 6

OpenStudy (anonymous):

yes i do so i write it out like this 6-y=6? or how would i write that out? idk why im so confused

OpenStudy (radar):

The sum of x pounda of peanuts plus y pounds of cashew is 6 pounds would be written out like this: x + y = 6

OpenStudy (radar):

That would be the first equation for the "system" of equations needed to solve this problem.

OpenStudy (anonymous):

okay cool and the second equation is 3.50x + 5.05y=$24.10 so then for part b how do i find that

OpenStudy (radar):

Yes, that shows the cost (money). Now you have two equations each with two unknowns. We must convert these two to one and here is how it is done; Take the first equation x + y = 6, and express x in terms of y getting s=6-y. all you are doing is subtracting y from both sides. A legal algebra move that you should be familiar with. Do you follow this far?

OpenStudy (radar):

Now you substitute 6-y for every x you see in the 2nd equation. I think I made a typo in my post above. x + y = 6, then x=6-y

OpenStudy (radar):

Here is what that would look like 3.50(6-y) + 5.05y =24.10 21 -3.50y + 5.05y = 24.10 1.55y = 24.10 - 21 1.55y = 3.10 y=2, or 2 pounds of cashews since x+ y =6 x must be 4, or 4 pounds of peanuts.

OpenStudy (anonymous):

wow i was way off but that makes wayyyyy more sense thank you so much!!

OpenStudy (radar):

Good luck with your studies.